OpenStudy (anonymous):

Write the equation of the line that has a slope of -2 and a y-intercept of -6. Write the equation of the line that passes through the point (6, -2) and has a slope of -2/3.

OpenStudy (anonymous):

y=-2x-6

OpenStudy (anonymous):

is that the first 1 or the second 1 durviel?

OpenStudy (anonymous):

Slope-intercept form is written \(\Large y = mx + c\) where m is the slope and c the y-intercept. If you know the slope and a point, you can work out what the intercept is by evaluating the slope-intercept equation and solving for c.

OpenStudy (anonymous):

first. i working on second

OpenStudy (anonymous):

\[\Large \begin{array}{l} y = mx + c\\ y = - \frac{2}{3}x + c\\ - 2 = - \frac{2}{3} \cdot 6 + c\\ - 2 + \frac{{12}}{3} = c\\ c = 2 \end{array}\]